OpenStudy (anonymous):

You have checked out b books from the library, where b is a number greater than 5. You plan to read three of the books and return them at the end of this week. In how many ways can you pick the books you will read first, second, and third?

OpenStudy (anonymous):

A. b3 ways B. (b3 - 3b2 + 2b) ways C. (b3 - 9b2 + 27b - 27) ways D. (5b - 3) ways

OpenStudy (kropot72):

There are b choices for the first book, (b - 1) choice for the second choice and (b - 2) choices for the third. \[\large Number\ of\ choices=b \times (b-1)\times (b-2)=you\ can\ multiply\ out\]

OpenStudy (anonymous):

so would it be D?

OpenStudy (arindameducationusc):

I guess Not D

OpenStudy (kropot72):

Partially multiplying out, the number of ways is given by\[\large b(b^{2}-3b+2)=you\ can\ multiply\ out\]

OpenStudy (anonymous):

Honestly, this makes no sence to me

OpenStudy (arindameducationusc):

okay, do you know distributive law?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I got B. Is that correct?

OpenStudy (kropot72):

Yes, b is the correct choice of answer.

OpenStudy (anonymous):

OKay, thank you

OpenStudy (kropot72):

You're welcome :)